War Log Explorer
TF REPEL 173 BSB
TF REPEL 173 BSB
Type of unit
Updated by group
292105Z TF REPEL CLP JAF TO BAF
Take care; definitions may be wrong.
Limited script-free view:
FROM: A/173rd BSB TO: BSB Battle Captain SUBJECT: TF REPEL CLP JAF to BAF Size and Composition of Patrol: 24 x US, 1 x TERPs A. Type of patrol: Mounted B. Task and Purpose of Patrol WILDCARD CLP conducts Convoy Logistics Patrol, between FOB FENTY and Bagram, IOT retro TF BAYONET units in N2KL.. C. Time of Return: 2105Z 29JUN07 D. Routes used and approximate times from point A to B: From Grid/FOB To Grid/FOB Route Travel JAF SP 1630z MSR Illinois ANP LU/RP 2 1630z 42S WD 28200 22500 ANP LU/RP 3 2000z 42S WD 25600 25800 BAF RP 2105z MSR Nevada Disposition of routes used: RTEs throughout our AO were green ATT. E. Enemy encountered: None F. Actions on Contact: N/A G. Casualties: N/A H. Enemy BDA: N/A I. BOS systems employed: N/A J. Final Disposition of friendly/enemy forces: N/A K. Equipment status: After Mission PMCS conducted upon arrival to motorpool L. M. Local Nationals encountered: LN# CP Name Village Tribe Approx age N/A N. Disposition of local security: None O. HCA Products Distributed: None P. PSYOP Products Distributed: None Q. Atmospherics: (reception of HCA, reactions to ANSF and Coalition forces, etc): None R. Reconstruction Projects QA/QC: N/A S. Afghan Conservation Corps nominations/Status: N/A T. Conclusion and Recommendation (Patrol Leader): Convoy brief 1500z TTPs, and safety brief delivered. CLP SP time 1630z. Initial part of the route was quiet, JAF ANP link went well. Travel through the mountains was clear and went without issues. We saw the same barrels that were reported on the last CLP at WD 4384 2363. The linkup with ANP in Kabul occurred successfully, and again movement through Kabul was quiet. There were no issues from the edge of Kabul to BAF. This evenings movement was all around quiet. After the recommendations, there are several grid coordinates that might be of interest/concern. Recommendations: There are several items being reported along the route that should be cleared, but are still present. Included after the recommendations are the grid coordinates of each location and the item of concern at that location. Additionally, have a BFT SME check out all BFTs. No reports seemed to be transmitted in either direction during the duration of the mission. Items of Concern: 42SXD 0268 1756 badly wrecked car on the side of the road 42SWD 9279 2432 additional piles of dirt and barrels 42SWD 8726 1888 large 1 meter deep hole in middle of road 42SWD 8610 1857 several cut sections in pavement of road 42SWD 8221 1800 van hit by ASV on side of road (STILL) 42SWD 7342 2516 another badly wrecked car o.s.o.r. 42SWD 6510 3005 piles of rocks/barrels o.s.o.r. 42SWD 5446 2711 more rocks/barrels o.s.o.r. 42SWD 4384 2363 still more barrels o.s.o.r. 42SWD 4011 2531 even more rocks and barrels o.s.o.r. Also- previously reported military trailer STILL present. Nothing Follows.